The Whitworth quick return motion mechanism has the driving crank AC=150 mm long.
The distance between fixed centres CD= 100 mm. The line of stroke of the ram passes
through the centre of rotation of the slotted lever whose free end is connected to the
ram by a connecting link. Find the ratio of time of cutting to time of return.
1
Expert's answer
2020-09-17T14:25:29-0400

Quick return ratio = {Time of cutting/time of return}

"\frac{a}{b}"

"360-\frac{a}{a}"

AC2 =150 mm ,CD = 100mm

Cos "\frac{b}{2}" = "\frac{CD}{AC}"

  = "\frac{100}{150}"

"\frac{b}{2}" = 48.2°

b = 96.4°

QRR = "\frac{360-96.4}{96.4}"

=2.735